### v1

# STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration

Abstract
Chiplet-based architectures and advanced packaging has emerged as transformative approaches in semiconductor design. While conventional physical design for 2.5D heterogeneous systems typically prioritizes wirelength reduction through tight chiplet packing, this strategy creates thermal bottlenecks and intensifies coefficient of thermal expansion (CTE) mismatches, compromising long-term reliability. Addressing these challenges requires holistic consideration of thermal performance, mechanical stress, and interconnect efficiency. We introduce STAMP-2.5D, the first automated floorplanning methodology that simultaneously optimizes these critical factors. Our approach employs finite element analysis to simulate temperature distributions and stress profiles across chiplet configurations while minimizing interconnect wirelength. Experimental results demonstrate that our thermal structural aware automated floorplanning approach reduces overall stress by 11% while maintaining excellent thermal performance with a negligible 0.5% temperature increase and simultaneously reducing total wirelength by 11% compared to temperature-only optimization. Additionally, we conduct an exploratory study on the effects of temperature gradients on structural integrity, providing crucial insights for reliability-conscious chiplet design. STAMP-2.5D establishes a robust platform for navigating critical trade-offs in advanced semiconductor packaging.
Index Terms:
Heterogeneous Integration, Automated Floor Planning, Chiplet Architecture, Thermal-Mechanical EvaluationI Introduction
The continuous advancement of semiconductor technology has driven an ever-increasing demand for higher performance, greater functionality, and enhanced integration density in electronic systems. Chiplet-based architectures, particularly 2.5D integration, have emerged as promising solutions in the electronics industry, bridging the gap between chip and package sizes by mounting multiple dies on a single package substrate. This approach enhances capacity and performance by facilitating heterogeneous integration of chiplets with varied functionalities and material properties into unified systems [1]. Advanced packaging technologies, including Through-Silicon-Via (TSV) fabrication methods and multi-level assembly techniques, have further solidified the reliability and affordability of 2.5D packaging [2].
Recent advancements in chiplet architectures and advanced packaging technologies have revolutionized system design, offering viable pathways to address the slowdown of Moore’s Law. Successful commercial implementations leverage chiplet architecture to integrate 7nm CPU dies with 12nm I/O dies for optimal performance and cost [3, 4]. Recent advances have further demonstrated scalability and efficiency in chiplet-based systems using heterogeneous 2.5D packaging techniques for workload-dependent configurations [5]. These innovations have catalyzed the development of diverse integration strategies, leveraging silicon interposers, bridge-chip technology, active interposers, and 3D interconnects [6, 1, 7]. Chiplet integration thus allows for scalable architectures, flexibility in node selection, and significant reuse of off-the-shelf intellectual properties (IPs)[8, 9]. However, substantial challenges persist, notably in inter-die communication, mixed-node integration, and thermal management[10, 11, 12].
Thermal management has become increasingly critical in 2.5D chiplet systems, driven by higher integration densities and power densities in advanced semiconductor nodes. Extensive research efforts employing finite element analysis and computational fluid dynamics have studied thermal behaviors and factors influencing package warpage, die stress, and solder joint reliability [13, 2, 6]. Nevertheless, while thermal considerations have been extensively explored, mechanical stress implications, particularly arising from the mismatch in the Coefficient of Thermal Expansion (CTE) among heterogeneous dies and interposers, remain inadequately investigated. Thermomechanical stress poses substantial reliability risks in 2.5D packages, warranting integrated consideration alongside thermal and interconnect optimization [14].
To address these challenges, recent methodologies have started exploring chiplet placement strategies that incorporate thermal considerations. One such approach, TAP-2.5D, strategically inserts spacing between chiplets to jointly optimize thermal performance and wirelength [15]. However, existing methodologies predominantly focus on thermal and interconnect objectives, neglecting a critical component—the mechanical stresses due to thermal expansion mismatches—which significantly affect long-term reliability.
In contrast, this paper presents STAMP-2.5D: Structural and Thermal Aware Methodology for Placement in 2.5D Integration, the first automated placement methodology that holistically integrates thermal performance, mechanical stress, and interconnect wirelength optimizations. Our proposed framework utilizes finite element analysis (FEA) simulations using ANSYS to concurrently evaluate thermal and mechanical stress profiles, enabling effective exploration of critical design trade-offs and ensuring robust reliability in advanced packaging.
II Literature Review
Effective thermal-mechanical stress management in 2.5D chiplet architectures remains a pivotal challenge, driving extensive research in both component-level analysis and system-level optimization. Early efforts, such as Jung et al. (2011), aimed at reducing computational complexity in reliability analysis by employing linear superposition of stress tensors to assess through-silicon vias (TSVs) stresses, which provided computational efficiency compared to traditional finite element methods (FEM)[16]. Nevertheless, this approach was limited to isolated TSV structures, neglecting integrated system-level implications and optimizations. Similarly thermal stresses in copper interwafer interconnects bonded with benzocyclobutene (BCB) are analysed offering insights into interwafer via stability yet without considering full-chip or system-wide interactions[17]. More recently, thermal-stress management by integrating design-time and run-time methodologies to minimize thermo-mechanical stresses in 3-D ICs have been ivestigated. Their work emphasized the importance of dynamic thermal gradients but lacked an automated optimization strategy explicitly targeting chiplet placement for simultaneous thermal and mechanical optimization [18].
Multi-scale and multi-fidelity modeling frameworks have emerged as promising techniques to bridge component-level and system-level analyses. Studies have integrated detailed thermal and mechanical analyses using finite element simulations with compact modeling methods, enabling performance predictions across multiple scales from individual components to complete systems [13, 19]. These approaches effectively reconcile localized stress evaluations with global thermal predictions, yet notably lack automated optimization algorithms necessary for practical placement decisions in chiplet-based systems.
Concurrently, automated placement methodologies have made significant strides in optimizing various design objectives, predominantly thermal performance, interconnect wirelength, and power efficiency. Prominent studies include those leveraging mathematical programming, heuristic-driven optimization, and machine-learning-based methods [20, 21, 15, 22, 23]. For instance, TAP-2.5D by Ma et al. (2023) explicitly introduced thermally-aware placement techniques, emphasizing insertion of spacing to reduce temperatures [15]. Chiou et al. (2023) and Deng et al. (2024) similarly utilized heuristic and reinforcement learning-based methods for thermal-aware floorplanning but did not explicitly incorporate mechanical stress as an objective or constraint in their methodologies [21, 23]. Although these studies represent significant advances in thermal and interconnect optimization, structural reliability and the critical implications of mechanical stress, especially arising from thermal gradients and interposer deformation, have been notably neglected. There is an critical need for automated placement methodology to jointly optimize chiplet configurations for thermal management, mechanical stress minimization, and interconnect wirelength efficiency.
III Methodology
This section describes the methodology for optimizing thermal, mechanical, and interconnect performance in 2.5D chiplet systems. We use a Simulated Annealing (SA) framework to balance thermal, mechanical, and interconnect objectives, with performance evaluated through ANSYS-based Finite Element Analysis (FEA) simulations. Mechanical reliability is assessed using von Mises stress, while wirelength is estimated to ensure efficient interconnect routing, enabling comprehensive optimization of chiplet-based architectures.
III-A Simulated Annealing
Our simulated annealing framework integrates a dynamically adaptive cost function that rigorously balances thermal, mechanical, and dimensional constraints. The process begins by calculating adaptive weights—denoted as , , and —which adjust based on the current and candidate design parameters, thereby allowing the algorithm to respond sensitively to variations in operating conditions.
Temperature Weight (): Considering 23∘C as the ambient baseline, the temperature weight is set to zero for conditions where both the old and new temperatures are below 75∘C. For higher temperatures, is defined as:
Here, the first term provides a base penalty starting at 0.1, which scales with the temperature’s deviation from the ambient (23∘C), capping at 0.5. The subsequent normalization factor ensures that the weight is appropriately scaled over the interval from 60∘C to 100∘C, capturing the increased significance of temperature beyond nominal operating conditions.
Stress Weight (): The weight for stress is computed to impose a progressively stronger penalty as the system approaches its mechanical limits. This is achieved by incorporating a nonlinear sensitivity term via an exponent:
represent the stress values of the current and candidate states, while is the maximum allowable stress. The nonlinearity introduced by the exponent of 1.5 ensures that as the stress approaches its upper limit, even marginal increases are penalized disproportionately, thereby safeguarding the system’s structural integrity.
Length Weight (): The length weight is derived as the residual factor needed to balance the overall weight distribution:
. This formulation accounts for the normalized change in length between iterations, ensuring that variations in dimension are factored into the cost function in a balanced manner relative to temperature and stress. The overall cost for a design is computed as a weighted sum of the normalized deviations in temperature, length, and stress. When normalization ranges are available, the cost functions for the current and candidate solutions are given by:
| old_cost | (4) | |||
| new_cost | ||||
If the normalization ranges are not applicable, a linear formulation is adopted instead. The difference in cost is defined as: . A positive (i.e., a decrease in cost) results in the immediate acceptance of the candidate solution. Conversely, if is negative, the candidate solution is accepted with a probability determined by the Boltzmann criterion: , where is the current annealing temperature. This probabilistic acceptance allows the algorithm to escape local minima by occasionally accepting solutions with higher costs, particularly during the early stages when is high. As the system cools, the likelihood of accepting suboptimal solutions diminishes, guiding the process toward convergence on a near-optimal configuration.
III-B Von Mises Stress
Mechanical reliability in 2.5D chiplet systems is typically assessed using von Mises stress, a scalar metric derived from the stress tensor components. Von Mises stress is particularly relevant for evaluating the likelihood of yielding or permanent deformation in materials subjected to complex loading conditions such as integrated circuits[18, 16]. Mathematically, von Mises stress is defined as:
where represents the von Mises stress, the terms , , and denote normal stresses in the respective coordinate directions, and , , and represent the shear stresses. Our approach employs this metric within ANSYS finite element analysis (FEA) to accurately capture structural stress profiles across heterogeneous chiplet arrangements.
III-C Wirelength estimation
Finding the minimum wirelength for inter-chiplet communication while satisfying bandwidth requirements is an NP-hard problem in multi-chiplet architectures. To accurately estimate interconnect wirelength, we adapted the technique from TAP-2.5D [15], integrating it into our multi-objective optimization framework. Our approach employs a Mixed-Integer Linear Programming (MILP) solver with an objective function that minimizes total wirelength:
where represents the distance between chiplets and , and denotes the number of wires connecting them.
The solver takes as input: (1) chiplet placement from our thermo-mechanically aware placement algorithm, (2) microbump resources for inter-chiplet communication, and (3) bandwidth constraints. Flow conservation constraints ensure the net flow from source to sink meets bandwidth requirements, while all intermediate nodes maintain zero net flow. The solver outputs the optimal routing and total wirelength with an average runtime of five seconds, enabling efficient design space exploration.
III-D Overview of Automated Floorplanning Algorithm
The proposed STAMP-2.5D methodology leverages a Simulated Annealing (SA) optimization framework to holistically optimize thermal management, structural reliability, and interconnect efficiency in chiplet-based systems. The algorithm initiates from an initial feasible chiplet placement—generated by adapting a widely-used automated floorplanning technique [24]—and iteratively explores improved configurations through stochastic perturbations.
At each iteration, the algorithm perturbs chiplet positions or orientations to generate candidate solutions, which are evaluated using an integrated cost function accounting for thermal performance, structural stress (von Mises), and interconnect wirelength. Thermal and mechanical stress profiles for candidate configurations are obtained through finite element analysis (FEA) simulations in ANSYS, ensuring accurate stress characterization. Interconnect wirelength is concurrently estimated using a MILP-based solver. Candidate configurations are then probabilistically accepted based on the Simulated Annealing criterion, enabling comprehensive exploration of the solution space and avoidance of local minima.
Fig. 1 provides a clear depiction of the iterative STAMP-2.5D floorplanning algorithm and highlights the integrated evaluation process.
III-E Modeling
| Components | Materials |
| Substrate | FR-4 | 0.3 | 13.0 | 20 | ||||||||
| C4 Bump | Tin-Lead (60-40) | 50 | 20.5 | 20 | ||||||||
| bump | SAC (SnAgCu) | 50 | 20 | 50 | ||||||||
| Interposer | Silicon Anisotropic | 148 | 2.6 | 150 | ||||||||
| Chiplets | Silicon Anisotropic | 150 | 3.1 | 130 | ||||||||
| TIM | Indium | 86 | 29 | 10 | ||||||||
| Heat Sink | Copper Alloy | 398 | 16 | 120 | ||||||||
| Underfill | SiO2 | 1.4 | 0.5 | 80 |
Our simulation framework employs ANSYS for coupled thermal-structural finite element analysis (FEA) of 2.5D chiplet systems, rigorously capturing multi-physics interactions across a detailed layer stack, as illustrated in Fig. 2. The stack comprises, from bottom to top, the substrate, C4 bumps, interposer, microbumps, chiplets, thermal interface material (TIM), and the heat sink. Each layer is modeled with its appropriate geometric and physical attributes. The bumps are approximated as cylinders and the heat sink is modeled as a parametric finned structure, with key parameters such as fin height, base thickness and pitch optimized to ensure efficient convective cooling while maintaining mechanical stability. Table I summarizes the materials and their properties for all components.
A refined mesh is applied in regions exhibiting steep thermal gradients and high stress concentrations—especially near interfaces and power-dense zones—to guarantee numerical precision. All interfaces (e.g., between chiplets and interposer, and between interposer and heat sink) are assumed to be perfectly bonded, thereby focusing the analysis on intrinsic material behavior while abstracting interfacial phenomena.
The simulation is conducted under steady-state conditions with a fixed ambient temperature of 23°C, and forced convective boundary conditions are applied on the heat sink surfaces to replicate realistic active cooling scenarios. By leveraging ANSYS’s multi-physics capabilities, our framework concurrently evaluates heat conduction and the resultant mechanical stresses. Specific experimental parameters, including localized power densities and variations in boundary conditions, are discussed in detail in the subsequent section.
III-F Experimental Setup
Our experimental evaluation utilized three representative architectures: Micro150, Ascend910, and MultiGPU, which cover various chiplet configurations found in research and industry applications, as illustrated in Fig. 3. We implemented these architectures on a 2.5D package with a silicon interposer of dimensions 45mm × 45mm for ascend910 and 50mm x 50mm for the rest and thickness 100m. Each architecture consists of multiple chiplets with varying dimensions and functionalities.
For our thermal-mechanical-aware simulated annealing-based floor planning experiments, we maintained a consistent chiplet thickness of 150 across all chiplet types. This standardized thickness was chosen to represent typical manufacturing practices for high-performance chiplets. However, in our exploratory studies discussed in the results section, we varied the chiplet thickness to investigate the impact on thermal and mechanical performance. For thermal simulation, we applied non-uniform power density distributions across chiplets detailed in Fig. 3.
The boundary conditions for our thermal simulations included an ambient temperature of 23°C with forced convective cooling applied to the heat sink. We used architecture-specific heat transfer coefficients to model different cooling scenarios: 580 W/m²·K for Ascend910, 720 W/m²·K for Micro150, and 950 W/m²·K for MultiGPU [25]. These values reflect the different thermal requirements of each architecture, with higher-power systems requiring more effective cooling solutions. The bottom surface of the package substrate was set to a natural convection boundary condition with a heat transfer coefficient of 10 W/m²·K.
We performed all finite element analyzes using ANSYS Mechanical for static thermal structural analyzes, with a medium mesh consisting of approximately forty thousand elements, with particular refinement near interfaces and areas with high thermal gradients. The mesh convergence study ensured that our results were independent of mesh resolution.
Our optimization algorithm parameters included an initial annealing temperature of 1.0, a cooling rate of 0.95, and 50 iterations per temperature level, with a total of 4,500 iterations for each optimization run. These parameters were empirically determined to balance exploration and exploitation in the design space.
IV Result and Analysis
We present experimental results from our thermal-mechanical aware floorplanning algorithm and analyze the critical relationship between temperature gradients and von Mises stresses in heterogeneous chiplet systems.
IV-A Thermal mechanical aware automated floor planning
We quantify temperature distributions, von Mises stress, and total interconnect wirelength through Ansys-based finite element analysis (FEA) simulations. Table II summarizes our results, clearly demonstrating the advantages of STAMP-2.5D compared to conventional thermal floor planning methodologies, while highlighting essential trade-offs among thermal management, structural reliability and connectivity efficiency.
Numerical results reveal compelling trade-offs. In the Micro150 architecture, our integrated approach notably reduces von Mises stress by 19.8%. This significant stress reduction is primarily driven by a symmetric chiplet placement strategy, which ensures uniform weight distribution across the interposer. Despite a modest temperature increase of only 2.97%, our method simultaneously achieves a wirelength reduction of 19.3%. The stress maps presented particularly in Fig. 4 clearly illustrate how symmetric placements lead to more balanced mechanical loads, contrasting sharply with the concentrated stress points typical of temperature-only optimization. This balanced weight distribution is critical in preventing localized solder joint failures and reducing interposer cracking during thermal cycling.
In parallel, the MultiGPU architecture demonstrates similar benefits. Here, our method achieves a 7.8% stress reduction, accompanied by a negligible temperature increase of just 0.39%, while improving the wirelength by 5.2%. The balanced placement proves particularly advantageous for managing the high power densities typical of GPU components within multi-GPU systems. Stress maps provided in Fig. 6 illustrate a clear reduction in localized stress near power-intensive GPU components. By more uniformly distributing thermal gradients across the interposer, our combined approach significantly reduces localized stress concentrations. This controlled management of thermal gradients effectively mitigates potential mechanical vulnerabilities caused by uneven expansion and contraction, as analyzed further in subsequent sections.
Consistent improvements are observed in the Ascend910 architecture, where our optimization strategy delivers dual advantages: a 4.5% reduction in von Mises stress alongside a 2.5% decrease in peak temperature. The temperature contour maps in Fig. 5 vividly demonstrate how structurally-aware placement effectively redistributes thermal concentrations around the critical Vitruvian component. Specifically, the sharp orange-red hotspot prominent in temperature-only optimization is transformed into smoother, more gradual temperature transitions with the combined method. This deliberate spreading of thermal gradients significantly enhances mechanical stability by minimizing localized expansion differentials and concurrently boosts overall thermal efficiency by utilizing a greater portion of the interposer’s surface area for effective heat dissipation. The STAMP-2.5D methodology thus represents not merely an optimal middle ground, but in most cases a superior solution: it avoids the high mechanical stress of temperature-only optimization, circumvents the elevated temperatures of stress-only approaches, and provides more reasonable wirelength than single-metric optimizations—ultimately delivering a balanced solution that simultaneously addresses thermal management, structural reliability, and interconnect efficiency in chiplet integration.
| Architecture | Cost Function | Temperature | Stress | Wirelength |
| (°C) | (MPa) | (mm) | ||
| Ascend910 | WT | 81.06 | 232.63 | 27200 |
| WS | 81.21 | 226.89 | 25895 | |
| WST | 79.04 | 222.14 | 27787 | |
| MultiGPU | WT | 85.69 | 280.82 | 82680 |
| WS | 87.07 | 275.85 | 98794 | |
| WST | 86.04 | 258.97 | 78358 | |
| Micro150 | WT | 95.24 | 291.02 | 104314 |
| WS | 99.04 | 224.00 | 93251 | |
| WST | 98.07 | 233.47 | 84227 |
IV-B Thermal Gradients and Structural Integrity Analysis
| Metrics | Ascend910 | Micro150 | Multi-GPU | ||||||
| WT | WS | WST | WT | WS | WST | WT | WS | WST | |
| Thermal Gradient Standard Deviation (Pa) | 466.11 | 495.40 | 410.59 | 932.96 | 1042.46 | 1050 | 708.47 | 711.46 | 653.01 |
| Hotspot Intensity Factor ((Tmax-Tavg)/Tavg) | 0.23 | 0.24 | 0.20 | 0.44 | 0.50 | 0.48 | 0.30 | 0.33 | 0.31 |
| Gradient-Stress Correlation Coefficient (r) | -0.25 | -0.22 | 0.02 | -0.25 | -0.15 | -0.15 | -0.22 | -0.24 | -0.24 |
| Stress Coefficient of Variance | 0.66 | 0.65 | 0.64 | 0.88 | 0.74 | 0.74 | 0.76 | 0.61 | 0.61 |
| Local Thermal-Mechanical Coupling Factor (Mean) (Pa.°C) | 59010.12 | 52912.96 | 33928.05 | 48198.98 | 39790.04 | 42687.99 | 34130.02 | 37143.14 | 40414.54 |
The peak stress that we observed in the previous subsection typically occurs on the interposer, highlighting the importance of analyzing how thermal gradients affect structural stability in our chiplet assemblies. Our analysis across three chiplet architectures reveals a fundamental insight: temperature gradients, rather than absolute temperatures, are the primary driver of mechanical stress in advanced chiplet assemblies. By incorporating mechanical stress considerations into our optimization framework, we achieve a sophisticated form of thermal-mechanical decoupling that enhances structural integrity without compromising performance.
The stress-aware optimization strategy consistently produces more uniform temperature gradient distributions across architectures, as illustrated in Fig. 7, which shows temperature gradient heat maps for all three optimization approaches. In Architecture 1 (Ascend910), this approach reduces the standard deviation of temperature gradients by nearly 12% and the mean gradient by 11.5%, creating more gradual thermal transitions that minimize stress concentrations. Architecture 3 (MultiGPU) shows similar improvements with an 8% reduction in gradient standard deviation. Remarkably, these improvements in gradient uniformity occur while maintaining nearly identical average temperatures around 65.6°C, confirming that the mechanical-thermal co-optimization methodology specifically targets harmful temperature discontinuities rather than merely reducing chip operating temperatures. The complete evaluation metrics for all three architectures are tabulated in Table III.
To quantify the degree of thermal-mechanical decoupling, we employ several specialized metrics. The Gradient-Stress Correlation Coefficient () measures the linear relationship between temperature gradients and mechanical stress, calculated as the Pearson correlation between and across all nodes in the finite element model. The Local Thermal-Mechanical Coupling Factor (LTMCF = ) quantifies how much stress is generated per unit of temperature gradient, with lower values indicating more effective decoupling of thermal effects from mechanical stress. The traditional Wirelength+Temperature optimization approach shows stronger negative Pearson correlations (around ) across all architectures, indicating that temperature gradients directly drive stress patterns. In contrast, the stress-informed floorplanning significantly weakens this correlation—shifting to near-zero (0.01) in Architecture 1 and reducing by over 40% in Architecture 2—demonstrating the ability to break the direct relationship between thermal gradients and mechanical stress.
Results from mechanical-aware optimization demonstrate promising stress reduction capabilities across architectures—with significant decreases of 19.8% in Micro150 and 7.8% in MultiGPU—despite maintaining or slightly increasing operating temperatures, indicating that temperature gradients likely play a more substantial role than absolute temperatures in thermal stress generation. Table III summarizes both thermal and stress metrics across all optimization strategies. These results clearly demonstrate that by explicitly considering mechanical stress during thermal optimization, chiplet designers can strategically manage temperature gradient distributions to minimize harmful stress concentrations, even in thermally challenging configurations. This mechanical-thermal co-optimization approach fundamentally transforms how thermal and mechanical behaviors interact in chiplet assemblies, offering a powerful strategy for enhancing long-term reliability in next-generation heterogeneous integration technologies.
V Conclusion and future work
In this paper, we introduced an automated floorplanning methodology that reduces thermally-induced structural stress in chiplet packages while optimizing wirelength with minimal temperature impact. Our results emphasize the importance of uniform temperature gradients for structural reliability. As future work, we aim to develop a multi-fidelity optimization framework to reduce runtime without compromising accuracy, and incorporate layer-specific CTE mismatch analysis within the metal stack. We also plan to explore interconnect technologies to further inform design strategies for robust chiplet integration.